OpenStudy (anonymous):

Find two positive numbers whose sum is 100 and the sum of whose squares is a minimum.

OpenStudy (anonymous):

Is it just me or is this problem pretty complex?

OpenStudy (anonymous):

this is what all of my problems are like... FML

OpenStudy (anonymous):

I don't understand squares is a minimum... hmmm

OpenStudy (anonymous):

x+y=100 x^2+y^2= something x^2 + (100-x)^2 = something differentiate set equal to zero

OpenStudy (anonymous):

i know how to do the problem now.. i just feel like these problems are really complex

OpenStudy (raden):

x+y=100 ----> y=100-x k=x^2+y^2 k=x^2+(100-x)^2 k=x^2+10,000-200x+x^2 k=2x^2-200x+10,000 u can use derivative like @Algebraic! said

OpenStudy (anonymous):

A rancher with 750ft of fencing wants to enclose a rectangular area and then divide it into four pens with fencing parallel to one side of the rectangle. a. Find a function that models the total area of the four pens. b. Find the largest possible total area of the four pens.

OpenStudy (raden):

the 2nd problem ?

OpenStudy (anonymous):

ya......

OpenStudy (raden):

 from the figure we can maka a models the total area of the four pens : 6x+4y=750 or 3x+2y=375 2y=375-3x y=(375-3x)/2 ...(1) the total area (A) = 4xy ....(2) subt (1) to (2) we get a function : A=4x(375-3x)/2 A=-6x^2+750x